Problem 3-12 Predetermined Overhead Rate; Disposing of Underapplied or Overapplied Overhead [LO3- 4) Luzadis Company makes fuRequired 1 Required 2 Required 3 Required 4 Compute the underapplied or overapplied overhead.Required 1 Required 2 Required 3 Required 4 Assume that the company closes any underapplied or overapplied overhead to Cost oRequired 1 Required 2 Required 3 Required 4 Assume that the company allocates any underapplied or overapplied overhead proporRequired 1 Required 2 Required 3 Required 4 How much higher or lower will net operating income be if the underapplied or over
Answer #1

1) Predetermine overhead rate = 900000/75000 = 12 per machine hour

Applied overhead = 60000*12 = 720000

Actual overhead = 850000

Under applied overhead = 130000

2) Journal entry

Date account and explanation Debit Credit
Cost of goods sold 130000
Manufacturing overhead 130000
(To record under applied)

3) Journal entry

Date account and explanation Debit Credit
Work in process 6500
Finished goods 32500
Cost of goods sold 91000
Manufacturing overhead 130000
(To record under applied)

4) Net income increase by (6500+32500) = 39000